Q5. Sketch the following signals. a) x[n] = 2X=o S[n – 2k] b) x(t) = Sto? 8(t – 2) dt
Answer #1

Q-5 K=3 in[n] = { [n-2k] Ko on putting the value of k n[n] = [n] + [n-2] + [n-4] + [n-6] psin) -2 -1 2 3 4 0 1 qon-2] 2 -1uine Y 10 10 10 } Am H-2) (b) x(t) = { $(E-2) dt as we know that ſo(t) dt = u(t) --- opci) in ep. ci) Put t = (2-2) idt = duUCA) nult-2)